A monopolist can produce two products of X and Y from a single production procedure with fixed proportion. (Q = QX = QY) The firm has marginal costs and revenues of MC = 12Q,

MRX = 54 – 6QX,

MRY = 126 – 12QY .

The producer should produce how many units?
Answer #1

Date Ans Marginal cost = 120 Marginal Revenue of x = 54 - 68x Marginal Revenue of Y = 126 - 12 Qy. we have equate Me and MR e